# AI Makes You Smarter, But None The Wiser: The Disconnect Between Performance and Metacognition

### v1

@@ 4.5.1. Analysis of prompts
4.5.2. Analysis of open-ended questions
@@
Table 3. Participant Approaches to AI Use
@@
High level of trust Participants relied heavily on AI (”blindly-trusted”), copying and pasting questions without critically assessing AI’s outputs or further inquiry. 58.94% (145 out of 246) 13.014 ± 3.062 16.844 ± 3.679
@@
5. Discussion
@@
This study offers insights into metacognitive monitoring in HAI by examining how users with varying competence interacted with AI during logical reasoning tasks. We explored the impact of AI on metacognitive accuracy, focusing on the DKE, user confidence, and AI literacy. Our findings reveal a significant inability to assess one’s performance accurately when using AI equally across our sample.
@@ 5.1. Effect of AI Literacy on Metacognition in Human-AI Interaction
While AI users in our sample outperformed those in Jansen et al. (2021), they consistently overestimated their performance by about four points, aligning with previous HCI research (Kosch et al., 2023; Villa et al., 2023; Kloft et al., 2024). The moderate correlation between estimated and actual scores (Table 2), with many participants rating themselves higher than the most skilled in the sample (5(a)) suggest that AI improves performance but leads to highly biased self-assessments.
This disconnect between actual and perceived performance mirrors earlier findings on overtrust and overreliance in AI systems (Kloft et al., 2024; Shekar et al., 2024; Lu and Yin, 2021). Overconfidence may impair users’ ability to evaluate their performance without AI, posing challenges for designing balanced human-AI interfaces. The classic DKE, where lower performers overestimate and higher performers underestimate their performance, disappeared with AI use, suggesting AI levels performance but does not correct inflated self-assessments. Metacognitive bias was doubled for the entire sample compared to Jansen et al. (2021).
We have also found an unexpected link between AI literacy and metacognitive accuracy. Participants with higher AI literacy were less accurate in self-assessments, contradicting the assumption that AI literacy improves metacognitive monitoring. Familiarity with AI may enhance the better-than-average effect (Brown, 1986; Zell et al., 2020), leading to overestimation of both relative and absolute performance.
@@
Qualitative data revealed varied perceptions of AI’s role, from a tool to a teammate, but these differences did not significantly affect performance or metacognitive accuracy, contradicting theories that suggest interaction framing impacts outcomes (Pataranutaporn et al., 2023; Villa et al., 2023). Regardless of user perception, the core metacognitive challenges in HAI persist.
@@ 5.2. Integration into theory
For HCI, high metacognitive bias leads users to overestimate their performance and over-rely on AI systems (Ma et al., 2024), reducing their ability to critically monitor HAI outcomes (Tankelevitch et al., 2024). When analyzing our results through Rasmussen’s SRK Model (Woods, 2009; Rasmussen, 1983), we find that heavy AI reliance shifted participants from reflective, knowledge-based processing to rule- or skill-based processing, reducing critical self-assessment and resulting in increased confidence but lower metacognitive accuracy. From a computational rationality perspective (Oulasvirta et al., 2022), this bias may be an adaptive response to AI presence, as participants may optimize perceived utility (e.g., efficiency) rather than monitoring HAI. This is supported by low metacognitive sensitivity and participants’ reliance on copy-pasting rather than higher-level metacognitive strategies (e.g., AI as collaborator). Such aligns with Villa et al. (2023), who found reduced error-processing when participants believed they were using sham AI. Therefore, our study provides further evidence of diminished metacognitive monitoring in HAI.
@@ 5.3. Limitations
While our study provides valuable insights into metacognitive monitoring in HAI, several limitations may affect the generalizability and interpretation of our findings. First, the apparent absence of the DKE, ”being unskilled and unaware” (Dunning, 2011), in our study may not fully reflect the underlying cognitive dynamics. AI interaction may have made participants equally skilled yet unaware of their performance rather than eliminating the DKE. AI-enhanced performance might have decoupled metacognitive judgments from cognitive performance (e.g., high confidence, low sensitivity, and high bias). Second, we relied on LSAT questions as our logical reasoning task. While these assess logical reasoning, they may not capture the diversity of real-world reasoning skills or domains. Furthermore, these tasks might overlap with the AI’s training data, limiting generalizability. However, since ChatGPT-4o’s performance was imperfect (68.25%), we believe our findings still apply to other tasks. Third, our focus on LSAT-based reasoning limits the scope of metacognitive biases across domains. Future research should use diverse tasks (e.g., writing creative texts with an LLM) to examine how AI interaction affects metacognitive monitoring and whether improvements in accuracy generalize across tasks. Lastly, we found little effect of different AI strategies on metacognitive accuracy and performance. The AI’s role, whether as a tool, collaborator, or teammate, did not significantly impact participants’ performance evaluation. Future research should explore how different AI roles (e.g., tool vs. collaborator) influence metacognitive accuracy and task performance.
@@ 5.4. Implications
Following Van Berkel and Hornbæk (2023), we identify three types of implications of our work: theoretical, design and methodological. For HCI theory, while biases in HAI have been studied (Liu et al., 2019; Kloft et al., 2024; Bertrand et al., 2022; Haliburton et al., 2024), little research has analyzed the metacognitive mechanisms underlying these biases and their impact on interaction (Tankelevitch et al., 2024). Applying metacognition theory to HAI (e.g., see Colombatto and Fleming (2023)), offers a new perspective for improving metacognitive monitoring in interaction design. Our findings also suggest that AI literacy alone is insufficient for optimal metacognitive abilities in HAI, challenging current political efforts like the EU AI Act (Parliament and of the European Union, 2024) that prioritize AI literacy for human-centered AI interaction. For design, we propose a new HAI interaction model that integrates metacognition and new design concepts. In the short term, designers should adjust interaction models, similar to Buçinca et al. (2021), and consider broader sociotechnical risks (Rafner et al., 2022). In the long term, HCI must develop a specific HAI interaction model that enhances metacognitive functions to support knowledge-based interactions (Rasmussen, 1983). Real-time sensing of metacognitive states (e.g., user agency) and adapting interfaces to cognitive states (Chiossi et al., 2023; Villa et al., 2024) could be crucial for optimizing HAI. Methodologically, our study introduces a new approach to analyzing biased decision-making in repeated AI interactions. Using computational methods, researchers can distinguish general biases from those emerging in post-decision processes (e.g., increased metacognitive noise), allowing for a more precise analysis of bias development in HAI.
@@ 6. Conclusion
Our study found that participants using an LLM to complete a logical reasoning task were unable to accurately assess their own performance, often overestimating it. With AI use, the DKE was eliminated. Based on these findings, we suggest developing new interfaces for interactive AI that are designed to enhance metacognition, allowing users to monitor their performance more accurately.

### v2

@@ 3.2.6. Analysis of prompts
Analysis of open-ended questions
@@
Table 3. Participant Approaches to AI Use in Study 1
@@
High level of trust Participants relied heavily on AI (”blindly trusted”), copying and pasting questions without critically assessing AI’s outputs or further inquiry. 58.94% (145 out of 246) 13.014 ± 3.062 16.844 ± 3.679
@@
3.3. Interim Dicussion
We found that using ChatGPT augmented our sample beyond a no-AI benchmark (i.e. AI augmentation) but that only a little more than half of our sample could surpass the AI alone (i.e. human-AI synergy). We found that most people overestimated their performance with AI and that there was no indication of a DKE when using the AI system. This may be due to participants’ tendency to not reflect on their performance and low metacognitive sensitivity, which is corroborated by our qualitative reports of people copying and pasting questions to the chat interfaces and then taking the AI’s answer without reflection.
4. Study 2: Incentivizing metacognitive thinking
To address the potential confound of a lack of motivation in our sample to engage in metacognitive monitoring, we conducted a second study in which participants received a monetary incentive for accurate judgments across the task (for a DKE study employing incentives, see Ehrlinger et al. (2008) Study 3). If participants monitor their performance when incentivized, the DKE could resurface. Given that Jansen et al. (2021) did not incentivize participants for accurate metacognitive judgments, this also mandated the sampling of a no-AI group within our study setup. We thus sampled another 250 participants for each the AI and the no-AI group to see if an incentive can motivate metacognitive monitoring and analyze the quantitative data. All data and analysis scripts for Study 2 can be found at https://osf.io/svax9/?view_only=cd176c08c2d4422baabf6bf815018bd7.
4.1. Method
We recruited 500 English-speaking participants located in the USA through Prolific. The sample was split into two groups: 250 participants completed the task without AI assistance (no-AI group) and 250 participants completed the task with AI assistance (AI group). Participants solved the same 20 logical reasoning problems used in Study 1. We did not collect the SNAIL for the no-AI group. Study 2 took, on average, 25 minutes to complete for the no-AI group and around 52 minutes for the AI group. Each participant was compensated 7 pounds per hour.
To motivate accurate self-assessment, participants in both groups were informed they would receive monetary compensation based on the accuracy of their performance estimates (compensation would be given to participants whose presumed number of correct answers closely matched their actual score). This incentive aimed to motivate participants to engage critically with the task and closely monitor their performance. All participants received full benefit compensation of 0.50 pounds regardless of their achieved performance. Similarly to Study 1, we included an attention check where participants were required to read a brief study and task description. They then answered two multiple-choice questions, one about the topic (logical reasoning) and another regarding how they could receive additional compensation (good judgment).
From 500 participants, we analyzed 452 participants (age: $M=37.24,SD=13.36$): 245 in the AI group and 207 in the no-AI group. Across both groups, 48 participants were excluded (3 for missing data, 3 for invalid performance estimates (exceeding 20 correct answers), and 42 for too low completion times.
202 participants identified as female, 242 as male, 6 as non-binary, 1 as two-spirit, and 1 who preferred not to disclose. Their highest educational degrees included 21 doctoral, 132 Master’s-level, 194 Bachelor’s-level, 64 upper secondary school, and 41 vocational qualifications. Regarding English proficiency, 391 participants identified as native speakers, 56 as fully fluent, 4 as conversationally fluent, and 1 as having basic proficiency.
For the AI group, a subset of participants (26) reported prior experience taking the LSAT. Their performance in this study ($M=13.50,SD=1.98$) was comparable to those without prior LSAT experience (n = 219; $M=13.25,SD=2.55$), and they were not excluded from the analysis. For the no-AI group, 25 participants reported prior LSAT experience, performing similarly ($M=9.50,SD=4.06$) to those without LSAT experience (n = 182; $M=9.52,SD=3.60$).
4.2. Results and Discussion
Participants in the AI group performed on average slightly worse as compared to AI alone ($M$ = 13.31, $SD$ = 2.44, $t(244)=-2.17,p=.031,d=-0.14$) but better than the no-AI group ($M$ = 9.71, $SD$ = 3.59, $t(450)=12.60,p<.001,d=1.18$). Therefore, we can assert that, on average, using AI has augmented performance but not that there is a synergy effect. In the AI group, 59.18% of participants scored higher than ChatGPT, with a total of 145 participants out of 245 surpassing its performance. In the no-AI group, 14.49% of participants scored higher than ChatGPT, corresponding to 30 participants out of 207, see also Figure 7. Therefore, performance in Study 2 mirrors Study 1.
Figure 7. Comparison of performance scores between the sample of participants interacting with ChatGPT and the sample of participants without AI. The blue density curve represents the distribution of performance scores in a sample of 245 participants using ChatGPT, showing a peak in performance at around 14 points. The yellow density curve corresponds to the sample of participants in the No-AI group, with lower overall performance scores. The vertical dashed line indicates the mean performance score in the ChatGPT simulation.
[…]
The graph compares performance scores between participants using ChatGPT and those without AI assistance from study 2. The ChatGPT group shows higher performance, peaking around 14 points, while the no-AI group has lower overall scores. The dashed line marks the mean performance in the ChatGPT simulation.
Investigating metacognitive accuracy for each sample, we find that in the AI group, participants overestimate their performance ($M$ = 17.13, $SD$ =3.16) which differed significantly from zero when subtracting individual performance ($t(244)=18.33,p<.001,d=1.17$). The same was found for the no-AI group ($M$ = 13.62, $SD$ = 4.14, $t(206)=11.81,p<.001,d=0.82$). Both quantile’s metacognitive accuracy differed from 0 for AI (Q1: $t(99)=-12.78,p<.001,d=-1.28$, Q4: $t(23)=-3.83,p<.001,d=-0.78$) as well as the no-AI group (Q1: $t(51)=-10.73,p<.001,d=-1.49$, Q4: $t(51)=-3.43,p=.001,d=-0.48$). Comparing estimates of estimated performance and performance of the first and the fourth quartile for each group, we find that the lowest quartile overestimates their performance relatively more when compared to the best-performing quartile (AI: $t(122)=4.06,p<.001,d=0.73$, no-AI: $t(102)=7.25,p<.001,d=1.42$), see 8(b). Therefore, the pattern of results in Study 2 regarding metacognitive accuracy also closely resembles Study 1. Comparing the confidence for correct and incorrect responses for the AI group, we find that, on average, participants are more confident for correct ($M$ = 85.95, $SD$ = 13.71) as compared to incorrect responses ($M$ = 82.57, $SD$ = 15.61;$t(244)=5.39,p<.001,d=0.34$). While confidence was descriptively lower for the no-AI group, we find the same pattern (correct: $M$ = 77.57, $SD$ = 14.85; incorrect: $M$ = 73.04, $SD$ = 16.73; $t(205)=6.33,p<.001,d=0.44$). A slight increase in confidence when accurate metacognition is incentivized is consistent with Ehrlinger et al. (2008).
Conducting a ROC analysis for each group, we found that most participants could distinguish between correct and incorrect answers. The mean AUC for the AI group ($M$ = .62, $SD$ =0.12) and the no-AI group ($M$ = .61, $SD$ = 0.11) differed from .5 (AI: $t(244)=15.30,p<.001,d=0.98$; no-AI: $t(205)=14.26,p<.001,d=0.99$ ), with most people exceeding the threshold of .5 AUC (AI: 196 of 245; no-AI: 172 of 207). The pattern of correlations of performance, metacognitive measures and AI literacy also resembled Study 1 in the AI group, see Table 4, and in the no-AI group, see Table 6.
Table 4. Correlation Table of Metacognitive measures and AI literacy as measured by the SNAIL in Study 2
$\Delta EP$
Estimate 0.71***
Performance -0.42*** 0.34***
$\Delta conf$0.05 -0.05 -0.01
$\mu conf$ 0.39*** 0.62*** 0.28*** -0.09
AUC -0.16* -0.09 0.09 0.49*** -0.23***
SNAIL TU 0.20** 0.18** -0.04 -0.26*** 0.21** -0.21***
SNAIL CA 0.14* 0.21*** 0.08 -0.08 0.14* -0.12 0.68***
SNAIL PA 0.09 0.19** 0.12 -0.13* 0.22*** -0.15* 0.75*** 0.83***
[…]
Applying our computational model using the same priors and sampler configuration as in Study 1, we find that both AI (Median = 0.63 (95% HDI [0.45, 0.85], $p_{b}$ = 0.0%) and no-AI (Median = 0.75 (95% HDI [0.58, 0.95], $p_{b}$ = 0.0%) show a metacognitive bias without distinguishing clearly between the AI group as compared to the non-AI sample; 18.1% of posterior samples were larger in the AI group as compared to the no-AI group, see also Figure 9A. Note that the discrepancy to Study 1, likely comes from the relatively lower precision, given the smaller sample size in Study 2. $\sigma_{k}$ indicating metacognitive noise was found to be above 1 for the no-AI group (Median = 1.53 (95% HDI [1.18, 1.96], $p_{b}$ = 0.1%), resembling the DKE pattern in Study 1, but centred around 1 for the AI group (Median = 1.13 (95% HDI [0.93, 1.36], $p_{b}$ = 10.1%), see also see also Figure 9B. 2.46% of posterior samples for sigma in the no-AI group exceed the AI group. Therefore, metacognitive noise does not scale the bias for the AI group, but it does for the no-AI group. We replicate the pattern of results in Study 1 again; see also Figure 9C. The difference in shape for Figure 9C and Figure 6C can be explained by the difference in range, especially regarding high performance, see Figure 3 and compare to Figure 7.
Overall, we can replicate the results of Study 1 in Study 2. Giving an incentive for accurate metacognitive judgements did not activate a DKE pattern for participants using AI. Nevertheless, we can see that the absolute levels of performance overestimation are slightly larger for the no-AI group in our sample (e.g., comparing estimated performance across studies).
(a) Scatter-plot of Estimated LSAT Scores as a function of Achieved LSAT Scores for the AI group in Study 2.
[…]
(b) Classical Dunning-Kruger Quartile-plot. Comparison of Mean Estimated and Mean Achieved Scores Across Quartiles for the AI group in Study 2.
[…]
Figure 8. Correlation of estimated and achieved LSAT score from different perspectives, individual 8(a) vs. quartile-level 8(b) for the AI group in Study 2.
[…]
Figure 9. Comparison posterior distributions with median and 95% HDI for the model parameters $b$ for bias in each group for bias (Plot A) and $\sigma$ (Plot B) for the second study. The posterior distributions of the AI group (in blue) and no AI group (in yellow). Plot C shows the average posterior predicted values for percent correct achieved (x-axis) and percent correct expected (y-axis) for each group. The s-shape around ideal metacognitive accuracy (grey line) indicates a DKE with low-performers overestimating their performance more than high-performers (yellow; no AI group).
[…]
Plot (a): comparison of posterior distributions for the bias parameter between the AI group and no-AI group (larger for AI group). Plot (b): comparison of posterior distributions of metacognitive noise for AI and non-AI groups, where the AI group shows less noise in judgement than the no-AI group. Plot (c): scatter plot showing the posterior predicted values for percent correct achieved (x-axis) and percent correct expected (y-axis) for each group. The s-shape around ideal metacognitive accuracy indicates a DKE with low-performers overestimating their performance more than high-performers.
5. General Discussion
@@
This paper offers insights into metacognitive monitoring in HAI by examining how users with varying competence interacted with AI during logical reasoning tasks. We explored the impact of AI on metacognitive accuracy, focusing on the DKE, user confidence, and AI literacy in two studies. Our findings reveal a significant inability to assess one’s performance accurately when using AI equally across our sample.
@@ 5.1. Effect of AI Literacy on Metacognition in Human-AI Interaction
While AI users in our sample outperformed those in Jansen et al. (2021), they consistently overestimated their performance by about four points, aligning with previous HCI research (Kosch et al., 2023; Villa et al., 2023; Kloft et al., 2024). The moderate correlation between estimated and actual scores (Table 2), with many participants estimating their joint performance with AI higher than the most skilled in the sample (5(a)), suggests that AI improves performance but leads to highly biased self-assessments.
This disconnect between actual and perceived performance mirrors earlier findings on overtrust and overreliance in AI systems (Kloft et al., 2024; Shekar et al., 2024; Lu and Yin, 2021). Overconfidence may impair users’ ability to evaluate their performance without AI, posing challenges for designing balanced human-AI interfaces. The classic DKE, where lower performers overestimate and higher performers underestimate their performance, disappeared with AI use, suggesting AI levels performance but does not correct inflated self-assessments. Metacognitive bias was doubled for the entire sample compared to Jansen et al. (2021) in Study 1, which was less pronounced in Study 2.
We have also found an unexpected link between AI literacy and metacognitive accuracy across both studies. Participants with higher AI literacy were less accurate in self-assessments, contradicting the assumption that AI literacy improves metacognitive monitoring. Familiarity with AI may enhance the better-than-average effect (Brown, 1986; Zell et al., 2020), leading to overestimation of both relative and absolute performance.
@@
Qualitative data revealed varied perceptions of AI’s role, from a tool to a teammate, but these differences did not descriptively show an effect on performance or metacognitive accuracy, see Table 3, contradicting theories that suggest interaction framing impacts outcomes (Pataranutaporn et al., 2023; Villa et al., 2023). Regardless of user perception, the core metacognitive challenges in HAI persist.
@@ 5.2. Integration into theory
For HCI, high metacognitive bias leads users to overestimate their performance and over-rely on AI systems (Ma et al., 2024), reducing their ability to critically monitor HAI outcomes (Tankelevitch et al., 2024). From a computational rationality perspective (Oulasvirta et al., 2022), this bias may be an adaptive response to AI presence, as participants may optimize perceived utility (e.g., efficiency) rather than monitoring HAI. This is supported by low metacognitive sensitivity and participants’ reliance on copy-pasting rather than higher-level metacognitive strategies (e.g., AI as collaborator). Such aligns with Villa et al. (2023), who found reduced error-processing when participants believed they were using sham AI. Therefore, our study provides further evidence of diminished metacognitive monitoring in HAI.
@@ 5.3. Limitations
While our study provides valuable insights into metacognitive monitoring in HAI, several limitations may affect the generalizability and interpretation of our findings. First, the apparent absence of the DKE (”being unskilled and unaware” (Dunning, 2011)), in our study may not fully reflect the underlying cognitive dynamics. AI interaction may have made participants equally skilled yet unaware of their performance rather than eliminating the DKE. AI-enhanced performance might have decoupled metacognitive judgments from cognitive performance (e.g., high confidence, low sensitivity, and high bias).
Second, we relied on LSAT questions as our logical reasoning task. While these assess logical reasoning, they may not capture the diversity of real-world reasoning skills or domains. Furthermore, these tasks might overlap with the AI’s training data, limiting generalizability. However, since ChatGPT-4o’s performance was imperfect (68.25%), we believe our findings still apply to other tasks.
Third, our focus on LSAT-based reasoning limits the scope of metacognitive biases across domains. Future research should use diverse tasks (e.g., writing creative texts with an LLM) to examine how AI interaction affects metacognitive monitoring and whether improvements in accuracy generalize across tasks. Fourth, we found little effect of different AI strategies on metacognitive accuracy and performance. The AI’s role, whether as a tool, collaborator, or teammate, did not impact participants’ performance evaluation (see again Table 3). Future research should explore how different AI roles (e.g., tool vs. collaborator) influence metacognitive accuracy and task performance.
Fifth, AI literacy was self-reported, which raises the same challenges inherent to the DKE: individuals with limited knowledge may overestimate their expertise. In line with this, self-reported AI literacy might not even directly translate into effective interactions with AI. Thus, how AI literacy links to performance and metacognition deserves to be further explored in future studies.
Sixth, participants were required to prompt ChatGPT at least once and then proceed with their decision, which may not reflect naturalistic usage patterns. In real-world scenarios, users might consult AI tools multiple times, ignore them entirely, or encounter more proactive AI systems offering suggestions unsolicited. Future studies should examine how varying degrees of user engagement with the AI and the AI’s proactiveness influence metacognitive accuracy, sensitivity, and bias.
Finally, exploring long-term learning scenarios—such as those examined in Bastani et al. (2024)—could illuminate how metacognitive processes evolve as individuals repeatedly interact with AI systems. Over extended periods, accurate metacognitive monitoring may become increasingly vital for achieving sustained performance gains (e.g., learning calculus with AI assistance and then taking an exam); our study cannot speak to the role of metacognition in AI-mediated learning contexts.
@@ 5.4. Implications
Following Van Berkel and Hornbæk (2023), we identify three types of implications of our work: theoretical, design and methodological. For HCI theory, while biases in human-AI interaction have been studied (Liu et al., 2019; Kloft et al., 2024; Bertrand et al., 2022; Haliburton et al., 2024), little research has analyzed the metacognitive mechanisms underlying these biases and their impact on interaction (Tankelevitch et al., 2024). Applying metacognition theory to HAI (e.g., see Colombatto and Fleming (2023)), offers a new perspective for improving metacognitive monitoring in interaction design. Our findings also suggest that AI literacy alone is insufficient for optimal metacognitive abilities in HAI. For design, we propose a new HAI interaction model that integrates metacognition and new design concepts. In the short term, designers should adjust interaction models, similar to Buçinca et al. (2021), and consider broader sociotechnical risks (Rafner et al., 2022). Furthermore, real-time sensing of metacognitive states (e.g., user agency) and adapting interfaces to cognitive states (Chiossi et al., 2023; Villa et al., 2024) could be crucial for optimizing HAI. Methodologically, our study introduces a new approach to analyze biased decision-making in repeated AI interactions. Using computational methods, researchers can distinguish general biases from those emerging in post-decision processes (e.g., increased metacognitive noise), allowing for a more precise analysis of bias development in HAI.
@@ 6. Conclusion
We found that participants using an LLM to complete a logical reasoning task were unable to accurately assess their own performance, often overestimating it. With AI use, the DKE was eliminated. Based on these findings, we suggest developing new interfaces for interactive AI that are designed to enhance metacognition, allowing users to monitor their performance more accurately.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Tables
Table 6. Correlation Table of Metacognitive Measures in Study 2 for the no-AI group
$\Delta EP$ Estimate Performance $\Delta conf$ $\mu conf$ AUC
$\Delta EP$
Estimate 0.68***
Performance -0.54*** 0.25***
$\Delta conf$0.22** -0.15* 0.11
$\mu conf$ 0.38*** 0.60*** 0.20** -0.18**
AUC -0.12 -0.13 0.01 0.57*** -0.24***
[…]
Appendix B Figures
(a) Scatter-plot of Estimated LSAT Scores as a function of Achieved LSAT Scores for the no-AI group in Study 2.
[…]
(b) Classical Dunning-Kruger Quartile-plot. Comparison of Mean Estimated and Mean Achieved Scores Across Quartiles for the no-AI group in Study 2.
[…]
Figure 10. Correlation of estimated and achieved LSAT score from different perspectives, individual 10(a) vs. quartile-level 10(b) for the no-AI group in Study 2.
[…]
Plot (a): scatter plot of estimated LSAT scores as a function of achieved LSAT scores across quartiles (Q1 to Q4) in the no-AI group in study 2. While some participants were accurate in their estimates, some participants were very wrong in estimating their performance accurately. Plot (b): quartile-level Dunning-Kruger plot for the no-AI group of study 2, showing that estimated scores are consistently higher than achieved scores across all quartiles.